# d2Cache: Accelerating Diffusion-Based LLMs via Dual Adaptive Caching

### v1

@@ 5.1 Experimental setup
Models, datasets, metrics and hardware. Following recent conventions (Wu et al., 2025), we evaluate our d2Cache on the Base and Instruct variants of two representative dLLMs (i.e., LLaDA-8B (Nie et al., 2025) and Dream-v0-7B (Ye et al., 2025)), which are denoted as LLaDA-Base/Inst and Dream-Base/Inst, respectively. Following Fast-dLLMWu et al. (2025), we evaluate our d2Cache on four benchmarks, including GSM8K (Cobbe et al., 2021), MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021), and Math-500 (Lightman et al., 2023), to assess performance across diverse reasoning and code generation tasks. The performance is reported in terms of task accuracy, which is evaluated using the lm-eval-harness framework (Gao et al., 2024). For fair comparisons, we report both inference throughput and latency, where throughput denotes the average number of tokens generated per second and latency denotes the average inference time per sample. All experiments are performed on NVIDIA 3090 24GB GPUs.
Baselines. We consider three baselines, including Vanilla and two representative approximate KV cache methods (i.e., dLLM-Cache Liu et al. (2025) and Fast-dLLM Wu et al. (2025)). For Vanilla, at each decoding step, the masked position with the highest confidence is replaced with its predicted token. For dLLM-Cache and Fast-dLLM, we employ their default configurations as reported in (Liu et al., 2025; Wu et al., 2025). More details are provided in appendix B of the Appendix.
Implementation details. Unless otherwise specified, the standard deviation $\sigma$ of the Gaussian function in Equation 3 is set to 10.0, the number of masked tokens selected per step is fixed at 32, the cumulative probability threshold $p$ is set to 0.1, and decoding is performed under the certainty prior.
5.2 Main results
The evaluation results on LLaDA-Inst and Dream-Inst are summarized in Table 1. Additional evaluation results on LLaDA-Base and Dream-Base can be found in Section C.1 of the Appendix (see Table 4). Notably, we observe that d2Cache achieves the best overall performance on average across four benchmarks, which delivers the highest throughput, the lowest latency, and the best score, consistently outperforming Vanilla, dLLM-Cache (Liu et al., 2025), and Fast-dLLM (Wu et al., 2025). Across all models and datasets, our d2Cache obtains an average 3.2×–4.0× speedup over Vanilla. Taking Dream-Inst on GSM8K as an example, our d2Cache improves the inference throughput from 2.62 to 12.25 tokens per second, leading to 4.7× inference speedup. More importantly, these substantial inference speedups are achieved without sacrificing accuracy, as the attainable score on average across four datasets remains comparable to or better than Vanilla. Furthermore, compared to recent representative approximate KV cache works (Wu et al., 2025; Liu et al., 2025), our d2Cache can also deliver better performance in terms of both inference efficiency and accuracy. For example, compared to Fast-dLLM, our d2Cache yields 1.5× inference speedup on Dream-Inst, while maintaining +3.2% accuracy on average across four datasets. These results clearly demonstrate the efficacy of d2Cache, which benefits from its two-stage fine-grained selection strategy.
@@ 5.2 Main results
NAR w/ Conf 57.5 3.0 42.1 26.4 32.7 51.6 34.2 26.8 3.2 29.0
NAR w/ Only CP 79.0 14.0 44.5 39.0 44.1 78.1 59.2 54.3 43.6 58.8
@@
5.3 Ablations and analysis
Certainty prior-guided decoding vs. confidence-based decoding. As discussed in Section 4.1, d2Cache naturally delivers an alternative decoding scheme: masked tokens can be decoded according to their certainty prior rather than their prediction confidence. To evaluate its efficacy, we further compare our certainty prior-guided decoding with the standard confidence-based decoding under the default non-autoregressive (NAR) setting. As shown in Table 2, our certainty prior-guided decoding delivers more reliable performance than the confidence-based decoding under the default NAR setting. In addition, when integrated into d2Cache, our certainty prior-guided decoding achieves the best average performance across four datasets, which clearly demonstrates its superiority.
@@ 6 Conclusion
Appendix A Relationships with concurrent works
@@
Appendix B Baseline hyperparameters
In this section, we provide more details about the hyperparameter configurations for the baseline methods (i.e., Fast-dLLM (Wu et al., 2025) and dLLM-Cache (Liu et al., 2025)) across different models and datasets. For Fast-dLLM, we closely follow common practices in prior work and set the block size to 32 for all models (Wu et al., 2025). For dLLM-Cache, we consider its key hyperparameters $K_{p}$ and $K_{r}$, where $K_{p}$ denotes the prompt refresh interval and $K_{r}$ denotes the response refresh interval. To ensure fair comparisons, we employ the default configurations as reported in (Liu et al., 2025), which are also summarized in Table 3. Furthermore, following (Nie et al., 2025), for the Instruct variants (i.e., LLaDA-Inst and Dream-Inst), all methods except d2Cache adopt block-wise semi-autoregressive (semi-AR) decoding with a block size of 32, whereas the Base variants (i.e., LLaDA-Base and Dream-Base) are evaluated in a fully non-autoregressive (NAR) manner.
@@ Appendix B Baseline hyperparameters
Appendix C Additional experimental results
C.1 Experimental results on the Base variants
In addition to the Instruct variants of LLaDA-8B (Nie et al., 2025) and Dream-v0-7B (Ye et al., 2025), we also conduct experiments on their Base variants, which are denoted as LLaDA-Base and Dream-Base, respectively. As shown in Table 4, our d2Cache consistently outperforms other approximate KV cache methods in terms of both average inference efficiency and accuracy across four datasets. Furthermore, we also note that the performance of Fast-dLLM (Wu et al., 2025) is substantially lower than that of the Vanilla baseline, particularly on Dream-Base with the MBPP dataset, where it exhibits a decline of about 20 points. This degradation aligns with prior findings that Base models are ill-suited for block-wise semi-autoregressive decoding (Nie et al., 2025). This further highlights the superiority of d2Cache over Fast-dLLM, as the latter heavily relies on block-wise semi-autoregressive decoding, which significantly restricts its applicability.
@@ C.1 Experimental results on the Base variants
MBPP
3-shot
Gen. Len. = 512
Vanilla 2.52 (1.0×) 195.59 39.2 2.81 (1.0×) 177.14 51.4
@@
HumanEval
0-shot
Gen. Len. = 512
Vanilla 5.02 (1.0×) 100.54 32.3 5.45 (1.0×) 92.11 51.2
@@
Math-500
4-shot
Gen. Len. = 256
Vanilla 3.14 (1.0×) 80.44 32.2 3.55 (1.0×) 71.54 39.0
@@
AVG
Vanilla 3.25 (1.0×) 122.24 43.5 3.62 (1.0×) 109.27 53.3
+ dLLM-Cache 8.28 (2.5×) 48.01 42.3 8.06 (2.2×) 52.61 50.4
+ Fast-dLLM 6.68 (2.1×) 62.59 41.9 7.03 (1.9×) 60.57 48.1
@@
C.2 More visualization results on attention rollout

### v2

@@ 5.1 Experimental setup
Models, datasets, metrics and hardware. Following recent conventions (Wu et al., 2025), we evaluate d2Cache on the Base and Instruct variants of two representative dLLMs (i.e., LLaDA-8B (Nie et al., 2025) and Dream-v0-7B (Ye et al., 2025)), which are denoted as LLaDA-Base/Inst and Dream-Base/Inst. Following dLLM-Cache (Liu et al., 2025), we evaluate d2Cache on six benchmarks, including GSM8K (Cobbe et al., 2021), MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021), Math-500 (Lightman et al., 2023), GPQA (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024) to assess performance across diverse reasoning, code generation and general tasks. The performance is reported in terms of task accuracy, which is evaluated using the lm-eval-harness framework (Gao et al., 2024). For fair comparisons, we report both inference throughput and latency, where throughput denotes the average number of tokens generated per second and latency denotes the average inference time per sample. All experiments are performed on NVIDIA 3090 24GB GPUs.
Baselines. We consider three baselines, including Vanilla and two representative approximate KV cache methods (i.e., dLLM-Cache and Fast-dLLM). For Vanilla, at each decoding step, the masked position with the highest confidence is replaced with its predicted token. For dLLM-Cache and Fast-dLLM, we employ their default configurations as reported in Liu et al. (2025); Wu et al. (2025). For Instruct variants, all baselines adopt block-wise semi-autoregressive decoding (semi-AR) with a block size of 32, whereas the Base variants are evaluated in fully non-autogressive (NAR) manner. More details are provided in appendix C of the Appendix.
Implementation details. Unless otherwise specified, the standard deviation $\sigma$ of the Gaussian function is set to 10.0, the number of masked tokens selected per step is fixed at 32, the cumulative probability threshold $p$ is set to 0.1, and the decoding is performed under the certainty prior.
@@
Semi-AR (Vanilla) 77.6 38.0 45.1 38.4 49.8 76.7 52.0 56.7 45.2 57.6
NAR w/ Conf 57.5 18.2 42.1 26.4 36.0 51.6 34.2 26.8 3.2 29.0
NAR w/ Only CP 79.0 38.8 44.5 39.0 50.3 78.1 59.2 54.3 43.6 58.8
Semi-AR w/ d2Cache 76.9 40.2 46.6 38.6 50.6 76.0 53.8 56.7 42.0 57.1
@@
5.2 Main results
The evaluation results on LLaDA-Inst and Dream-Inst are summarized in Table 1. Notably, we observe that d2Cache achieves the best overall performance on average across all benchmarks, which delivers the highest throughput, the lowest latency, and the best score, consistently outperforming all baselines. Across all models and datasets, our d2Cache obtains an average 3.5× speedup over Vanilla. Taking Dream-Inst on GSM8K as an example, our d2Cache improves the inference throughput from 2.62 to 12.25 tokens per second, leading to 4.7× inference speedup. More importantly, these substantial inference speedups are achieved without sacrificing accuracy, as the attainable score on average across six datasets remains comparable to or better than Vanilla. Furthermore, compared to dLLM-Cache and Fast-dLLM, our d2Cache can also deliver better performance in terms of both inference efficiency and accuracy. For example, compared to Fast-dLLM, our d2Cache yields 1.3× inference speedup on Dream-Inst, while maintaining +1.7% accuracy on average across six datasets. These results clearly demonstrate the efficacy of d2Cache, which benefits from its two-stage fine-grained selection strategy.
@@ 5.2 Main results
5.3 Ablations and analysis
Certainty prior-guided decoding vs. confidence-based decoding. As discussed in Section 4.1, d2Cache naturally delivers an alternative decoding scheme: masked tokens can be decoded according to their certainty prior rather than their prediction confidence. To evaluate its efficacy, we further compare our certainty prior-guided decoding with the standard confidence-based decoding under the default NAR setting. As shown in Table 2, our certainty prior-guided decoding delivers more reliable performance than the confidence-based decoding under the default NAR setting. We also observe that certainty prior–guided decoding and semi-AR decoding achieve comparable performance, because both approaches constrain the model to decode in a quasi left-to-right manner. Although they share a similar intuition, only the combination of certainty prior–guided decoding and d2Cache delivers the best performance among all evaluated configurations.
Figure 6: Hyperparameter sensitivity analysis of $p$, $k$, and $\sigma$ on Dream-Inst and HumanEval.
@@ 5.3 Ablations and analysis
Hyperparameter sensitivity analysis. To determine the optimal hyperparameters, we conduct systematic experiments on Dream-Inst and HumanEval with generation lengths of 256, 512 and 1024. As shown in Figure 6 (a-c), the number of masked tokens updated per step is the dominant factor: performance improves as $k$ increases but saturates—and may slightly decline—beyond $k=32$, indicating that $k=32$ offers the most stable gains across settings of $p$ and sequence lengths. The cumulative probability threshold $p$, which regulates the retained probability mass and thus affects throughput, does not monotonically improve performance with larger values. We additionally examine the Gaussian standard deviation $\sigma$ in Equation 3, which governs the locality of certainty-prior selection.
@@ 6 Conclusion
Appendix A The use of large language models
In this work, we employ large language models (LLMs) as general-purpose auxiliary tools, which are mainly used in the following two scenarios:
  • Writing and editing: LLMs assist in revising the manuscript by enhancing its clarity, grammar, and stylistic consistency.
  • Code generation: LLMs assist in programming tasks, including debugging and generating illustrative code snippets.
The authors are fully responsible for the entire content of this paper, including sections in which LLMs provide writing assistance. We note that LLMs are not involved in research ideation, experimental design, or data analysis, and therefore do not meet the criteria for authorship.
Appendix B Relationships with concurrent works
@@
Appendix C Baseline hyperparameters
In this section, we provide more details about the hyperparameter configurations for the baseline methods (i.e., Fast-dLLM (Wu et al., 2025) and dLLM-Cache (Liu et al., 2025)) across different models and datasets. For Fast-dLLM, we closely follow common practices in prior work and set the block size to 32 for all models (Wu et al., 2025). For dLLM-Cache, we consider its key hyperparameters $K_{p}$ and $K_{r}$, where $K_{p}$ denotes the prompt refresh interval and $K_{r}$ denotes the response refresh interval. To ensure fair comparisons, we employ the default configurations as reported in Liu et al. (2025), which are also summarized in Table 3.
@@ Appendix C Baseline hyperparameters
GPQA LLaDA-8B-Instruct 50 6
Dream-v0-7B-Base 100 8
Dream-v0-7B-Instruct 10 8
MMLU-Pro LLaDA-8B-Base 100 6
LLaDA-8B-Instruct 50 3
Dream-v0-7B-Base 25 2
Dream-v0-7B-Instruct 5 1
Appendix D Discussions
D.1 Memory overhead of caching
We conduct a thorough analysis and profiling of the memory overhead of caching. Note that the KV cache used by dLLMs consumes the same amount of memory as that required by an autoregressive LLM (ARM) of the same scale. Specifically, for sequence length $L$, number of layers $N$, and hidden dimension $d$, an ARM or a dLLM stores $2\times L\times N\times d$ floating-point values for the KV cache. d2Cache additionally stores an attention-rollout matrix of size $L\times L$, which is typically negligible. We report the peak memory usage on Dream Inst for a generation length of 1024 across four datasets. As shown in the Table 7, for example on GSM8K, where the average prompt length is approximately 800—resulting in a sequence length of roughly 1.8k—the additional memory consumption of d2Cache is nearly identical to that of Fast-dLLM (Wu et al., 2025).
D.2 Decoding order of dLLMs
In Section 4.1, we proposed certainty-prior–guided decoding, which forces the model to generate in a quasi–left-to-right order. A natural question arises: if dLLMs behave more like autoregressive models (ARMs), does this violate the original intention of enabling parallel, any-order generation? Here, we argue that the answer is not simply “no”.
Unlike ARMs, which can generate tokens only at the immediately adjacent next position, dLLMs produce predictions over the entire sequence, which is the source of their any-order generation capability. However, high-quality predictions are not available at all positions. Thus, selecting which tokens to decode—that is, determining the decoding order—is crucial for generation quality. As shown in Nie et al. (2025), LLaDA-Instruct tends to become prematurely overconfident in EOS tokens near the end of the sequence, and therefore proposes block-wise semi-autoregressive decoding (semi-AR), which constrains the model to decode from left to right at the block level while generating in parallel within each block. Compared with fully non-autoregressive decoding, block-wise semi-AR preserves the model’s sequential reasoning ability to a large extent, as shown in Table 2.
In our paper, experiments in Section 3.2 show that a dLLM consistently prefers to decode tokens close to known positions. This observation explains why block-wise semi-AR is effective: enforcing quasi-left-to-right generation ensures that each token is decoded only when the contextual information is sufficiently rich. Our certainty-prior decoding shares the same intuition, but provides a more conceptual formulation.
Although dLLMs need to decode in a quasi–left-to-right order to maintain sequential reasoning ability, they still retain substantially greater flexibility during generation. For example, when the model encounters a position where all next-token candidates have low confidence, an AR model must commit to one choice. In contrast, a dLLM can decode further positions and delay the decision until the extended context provides adequate evidence, thereby exploiting its non-AR modeling capacity. A concrete example is pronoun resolution in ambiguous contexts. Suppose the prompt is: “Alice thanked Mary because ____ had helped with the project”.
At the blank position, an AR model must immediately choose between “she” and “Alice”, even though the correct antecedent remains unclear without additional context. A dLLM, however, can tentatively consider both possibilities, continue decoding subsequent positions, and use the extended context to determine whether the sentence is likely to continue as “… she had provided key data”, or “… Alice needed assistance”, before committing to the final token.
Moreover, even under quasi-AR decoding, each masked token still attends to the entire context (unlike ARMs, where tokens can only attend to previous positions), so the original advantages of dLLMs, such as bidirectional modeling and parallel decoding, remain preserved.
Table 4: Comparisons of using only the first 5 layers to compute attention rollout (Rollout-5) and using all layers to compute attention rollout (Full-rollout) on Dream-Inst. Bold numbers indicate the best scores, and green texts denote the speedup ratios relative to the Vanilla method.
Dataset Method Throughput ↑ Score ↑
GSM8K 4-shot Gen. Len. = 256 Vanilla 2.622.62(1.0×) $76.7$
Rollout-5 12.6112.61(4.8×) $71.8$
[…]
MBPP 3-shot Gen. Len. = 512 Vanilla 2.732.73(1.0×) $52.0$
Rollout-5 13.1013.10(4.8×) $57.2$
[…]
HumanEval 0-shot Gen. Len. = 512 Vanilla 4.394.39(1.0×) $56.7$
Rollout-5 14.2014.20(3.2×) 62.2
[…]
Math-500 4-shot Gen. Len. = 256 Vanilla 3.513.51(1.0×) 45.2
Rollout-5 13.9913.99(4.0×) $40.2$
[…]
@@ D.2 Decoding order of dLLMs
MBPP 3-shot Gen. Len. = 512 Vanilla 2.52 (1.0×) 195.59 39.2 2.81 (1.0×) 177.14 51.4
@@
HumanEval 0-shot Gen. Len. = 512 Vanilla 5.02 (1.0×) 100.54 32.3 5.45 (1.0×) 92.11 51.2
@@
Math-500 4-shot Gen. Len. = 256 Vanilla 3.14 (1.0×) 80.44 32.2 3.55 (1.0×) 71.54 39.0
@@
GPQA 0-shot Gen. Len. = 256 Vanilla 6.27 (1.0×) 42.35 30.4 6.54 (1.0×) 40.55 32.8
+ dLLM-Cache 11.32 (1.8×) 22.69 31.0 11.12 (1.7×) 23.10 34.6
+ Fast-dLLM 12.67 (2.0×) 20.24 31.0 11.92 (1.8×) 21.45 31.5
@@
MMLU-Pro 5-shot Gen. Len. = 256 Vanilla 1.53 (1.0×) 143.45 38.1 2.13 (1.0×) 127.08 46.1
+ dLLM-Cache 6.86 (4.5×) 37.96 37.4 7.45 (3.5×) 34.81 44.6
+ Fast-dLLM 8.96 (5.9×) 28.83 40.0 9.42 (4.4×) 27.31 45.9
[…]
AVG Vanilla 3.47 (1.0×) 112.46 40.4 3.86 (1.0×) 100.79 48.7
+ dLLM-Cache 8.55 (2.5×) 42.12 39.6 8.46 (2.2×) 44.72 46.8
+ Fast-dLLM 9.61 (2.8×) 39.07 39.6 9.50 (2.5×) 43.09 49.1
[…]
D.3 Limitations and future work
Although d2Cache delivers substantial inference speedups across multiple models and datasets while maintaining comparable performance, several limitations have also emerged. Below we summarize these limitations and further outline potential directions for future work.
Larger-scale dLLMs. In this work, we closely follow recent representative practices (Wu et al., 2025; Liu et al., 2025) to evaluate d2Cache on LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025). We note that, at this moment, LLaDA-8B and Dream-7B are the only publicly available dense dLLMs. As future dLLMs continue to scale up in depth, width, and context length, their bidirectional attention patterns will become even more costly to maintain during decoding. This trend further highlights the importance of more effective caching schemes. We view extending d2Cache to larger-scale dLLMs—together with more effective caching schemes—as a promising direction for future work, especially as model sizes and application demands continue to explode.
Lightweight variants of attention rollout. Although attention rollout is not a performance bottleneck in our d2Cache, its cost can become significant when it is applied to larger-scale models. More efficient approximations are therefore desirable. We evaluate a lightweight variant that computes rollout using only the first five layers on Dream-Inst and four datasets. As shown in Table 4, reducing the rollout depth from 28 to 5 yields a slight improvement in decoding speed while noticeably degrading performance on math reasoning tasks (GSM8K and Math-500); in contrast, code-generation tasks (HumanEval and MBPP) exhibit minimal performance loss. Designing lightweight rollout variants that can identify key tokens still remains an important direction for future work.
Alternative scoring functions for contextual contribution. We currently employ a Gaussian function to characterize how a masked token influences its surrounding context. While this approach performs well empirically, more context-adaptive formulations may further enhance performance.
Appendix E Additional experimental results
Table 6: Comprehensive evaluation results on LLaDA-Inst (Nie et al., 2025) and Dream-Inst (Ye et al., 2025) with semi-AR parallel decoding. Bold numbers indicate the best results and green texts denote the speedup ratios.
Dataset Method LLaDA-Inst Dream-Inst
Throughput ↑ Latency(s) ↓ Score ↑ Throughput ↑ Latency(s) ↓ Score ↑
GSM8K 4-shot Gen. Len. = 256 Vanilla 2.77 (1.0×) 110.26 77.6 2.62 (1.0×) 85.94 76.7
Parallel 8.53 (3.1×) 33.95 77.6 13.93 (5.3×) 21.68 74.2
+ dLLM-Cache 25.62 (9.2×) 10.26 77.0 36.00 (13.7×) 8.21 74.3
+ Fast-dLLM 25.15 (9.1×) 10.65 77.6 32.75 (12.5×) 8.35 74.1
[…]
MBPP 3-shot Gen. Len. = 512 Vanilla 2.48 (1.0×) 199.90 14.4 2.73 (1.0×) 182.78 52.0
Parallel 29.92 (12.1×) 19.61 38.8 38.06 (13.9×) 14.95 51.6
+ dLLM-Cache 72.52 (29.2×) 7.94 38.0 89.31 (32.7×) 6.11 51.8
+ Fast-dLLM 43.29 (17.5×) 12.44 38.4 51.04 (18.7×) 10.40 52.4
[…]
HumanEval 0-shot Gen. Len. = 512 Vanilla 4.99 (1.0×) 105.76 45.1 4.39 (1.0×) 114.86 56.7
Parallel 15.74 (3.2×) 37.63 45.1 39.78 (9.1×) 20.53 51.8
+ dLLM-Cache 27.88 (5.6×) 20.54 48.2 48.50 (11.0×) 14.75 53.7
+ Fast-dLLM 25.14 (5.0×) 21.76 43.3 48.94 (11.1×) 13.67 57.3
[…]
Math-500 4-shot Gen. Len. = 256 Vanilla 3.08 (1.0×) 82.51 38.4 3.51 (1.0×) 71.05 45.2
Parallel 8.80 (2.9×) 31.99 38.4 12.75 (3.6×) 23.99 44.4
+ dLLM-Cache 17.90 (5.8×) 15.53 37.8 25.16 (7.2×) 12.06 43.0
+ Fast-dLLM 24.49 (8.0×) 11.01 37.4 28.68 (8.2×) 9.73 43.4
[…]
GPQA 0-shot Gen. Len. = 256 Vanilla 6.14 (1.0×) 43.34 25.2 6.43 (1.0×) 41.14 30.1
Parallel 44.61 (7.3×) 15.98 25.7 128.85 (20.0×) 2.87 31.3
+ dLLM-Cache 50.17 (8.2×) 10.78 28.4 174.83 (27.2×) 2.06 33.3
+ Fast-dLLM 42.66 (6.9×) 10.13 25.9 136.28 (21.2×) 2.31 34.6
[…]
MMLU-Pro 5-shot Gen. Len. = 256 Vanilla 1.76 (1.0×) 152.62 37.5 2.15 (1.0×) 126.31 47.9
Parallel 8.82 (5.0×) 58.86 37.2 17.87 (8.3×) 27.12 47.8
+ dLLM-Cache 19.76 (11.2×) 18.95 35.2 41.83 (19.5×) 8.88 48.9
+ Fast-dLLM 19.46 (11.1×) 16.07 37.1 35.46 (16.5×) 8.94 47.1
[…]
AVG Vanilla 3.54 (1.0×) 115.73 39.7 3.64 (1.0×) 103.68 51.4
Parallel 19.40 (5.5×) 33.00 43.8 41.87 (11.5×) 18.52 50.2
+ dLLM-Cache 35.64 (10.1×) 14.00 44.1 69.27 (19.0×) 8.68 50.8
+ Fast-dLLM 30.03 (8.5×) 13.68 43.3 55.53 (15.3×) 8.90 51.5
[…]
E.1 Experimental results on the Base variants
In addition to the Instruct variants of LLaDA-8B (Nie et al., 2025) and Dream-v0-7B (Ye et al., 2025), we also conduct experiments on their Base variants, which are denoted as LLaDA-Base and Dream-Base, respectively. As shown in Table 5, our d2Cache consistently outperforms other approximate KV cache methods in terms of both average inference efficiency and accuracy across six datasets.
Table 7: Performance comparison on Dream-Inst with a generation length of 1024.
Dataset Method Throughput (tokens/s) ↑ Latency(s) ↓ Score ↑ Memory (GB) ↓
GSM8K 4-shot Gen. Len. = 1024 Vanilla 1.54 (1.0×) 671.35 68.46 19.26
Fast dLLM 4.18 (2.7×) 245.29 67.85 19.39
dLLM-Cache 3.33 (2.2×) 308.62 68.76 20.28
[…]
Math-500 4-shot Gen. Len. = 1024 Vanilla 1.89 (1.0×) 541.04 43.6 19.15
Fast dLLM 4.43 (2.3×) 231.45 42.6 19.27
dLLM-Cache 2.75 (1.5×) 373.1 40.4 20.2
[…]
HumanEval 0-shot Gen. Len. = 1024 Vanilla 2.62 (1.0×) 393.47 56.71 19.06
Fast dLLM 4.77 (1.8×) 214.5 58.53 19.16
dLLM-Cache 3.03 (1.2×) 338.21 60.97 19.79
[…]
MBPP 3-shot Gen. Len. = 1024 Vanilla 1.95 (1.0×) 526.86 52.8 19.12
Fast dLLM 4.45 (2.3×) 229.91 52.4 19.23
dLLM-Cache 4.59 (2.4×) 223.05 54.2 19.94
[…]
AVG Vanilla 2.00 (1.0×) 533.18 55.39 19.15
Fast dLLM 4.46 (2.2×) 230.29 55.35 19.26
dLLM-Cache 3.43 (1.7×) 310.75 56.08 20.05
[…]